# A quantum Monte Carlo algorithm for arbitrary high-spin Hamiltonians

### v1

@@ V.3 Incorporation of bosons
As for the inclusion of bosonic degrees of freedom, we will see in this section that the creation and annihilation operators in the second-quantized basis correspond to permutation operators of infinite order (see also [19]).
@@
where $a_{i}$ are unknown integers. Equation (41) can be addressed using standard methods for Diophantine equations (see Appendix B).
@@ VI Summary and discussion
We presented a universal, parameter-free, Trotter-error-free quantum Monte Carlo scheme capable of simulating, for the first time, arbitrary high-spin Hamiltonians. We have demonstrated that the permutation matrix representation of Hamiltonians allows one to automatically produce QMC updates that are provably ergodic and satisfy detailed balance, thereby ensuring the convergence of the Markov chain to the equilibrium.
Our algorithm therefore enables one to study the equilibrium properties of essentially any conceivable high-spin system using a single piece of code. We illustrated this ability by producing results for the quantum Heisenberg model for two types of spin particles, namely, spin-$1$ and spin-$3/2$.
We have further shown that the methodologies devised here can be extended to other particle species and mixtures thereof.
We hope that the feasibility of such calculations and applicability of our code to a wide range of high-spin systems will contribute to the broader understanding of quantum magnetism and phase transitions in low-dimensional quantum systems.
We note that while our approach guarantees a correct equilibrium distribution of the Markov chain, the algorithm does not guarantee a universal rapid mixing of the Markov chain, nor does it resolve or aim to resolve the sign problem.
@@ Appendix A Cycle completion with gaps
Appendix B Solving systems of modular equivalences
A system of modular equivalences Eq. (35) can be rewritten as a system of Diophantine equations
where $k_{1},k_{2},\dots,k_{N}$ are additional unknown integers. The Diophantine system (45) can be solved by employing Hermite normal form of the corresponding matrix. For details, the reader is referred to Refs. [21, 22].

### v2

@@ V.3 Incorporation of bosons
As for the inclusion of bosonic degrees of freedom, we will see in this section that the creation and annihilation operators in the second-quantized basis correspond to permutation operators of infinite order (see also Ref. [35]).
@@
where $a_{i}$ are unknown integers. Equation (44) can be addressed using standard methods for Diophantine equations.
@@ VI Summary and discussion
We presented a universal Trotter-error-free quantum Monte Carlo scheme capable of simulating, for the first time, arbitrary high-spin Hamiltonians. We have demonstrated that the permutation matrix representation of Hamiltonians allows one to automatically produce QMC updates that are provably ergodic and satisfy detailed balance, thereby ensuring the convergence of the Markov chain to the equilibrium.
Our algorithm therefore enables one to study the equilibrium properties of essentially any conceivable high-spin system using a single piece of code. While our approach guarantees a correct equilibrium distribution of the Markov chain, the algorithm does not guarantee a universal rapid mixing of the Markov chain, nor does it resolve or aim to resolve the sign problem.
We illustrated this ability by producing results for the quantum Heisenberg model for two types of spin particles, namely, spin-$1$ and spin-$3/2$. To demonstrate the versatility of our approach, we have studied in addition the equilibrium properties of randomly generated Hamiltonians, which existing QMC techniques cannot simulate.
We have further shown that the methodologies devised here can be extended to other particle species and mixtures thereof. We hope that the feasibility of such calculations and applicability of our code to a wide range of high-spin systems will contribute to the broader understanding of quantum magnetism and phase transitions in low-dimensional quantum systems.